OpenStudy (anonymous):

write the following L0G form in exponential form: lne^-2=b. My online hw website I am using is saying it's incorrect when I put in 2^b=e.

OpenStudy (anonymous):

\[\ln(e^x)=x\]

OpenStudy (anonymous):

so if \[\ln(e^{-2})=b\] that makes \[-2=b\]

OpenStudy (anonymous):

it still says its incorrect :/ quite weird if you seem correct though